40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Mar 28 03:37:07 2013, Zach, Optics, Configuration, Test setup input optics progress OMC_test_setup_IO.jpg2013-03-27_05.48.20.jpg2013-03-28_02.29.36.jpg
    Reply  Fri Mar 29 08:55:00 2013, Zach, Optics, Configuration, Beam launched into fiber 
       Reply  Thu Apr 4 23:44:52 2013, Koji, Optics, Configuration, Beam launched into fiber 
          Reply  Fri Apr 5 18:18:36 2013, Zach, Optics, Configuration, AOM probably broken Isomet_1250C_zin_4_5_13.png
Message ID: 86     Entry time: Thu Mar 28 03:37:07 2013     Reply to this: 87
Author: Zach 
Type: Optics 
Category: Configuration 
Subject: Test setup input optics progress 

[Lisa, Zach]

Last night (Tuesday), I finished setting up and aligning most of the input optics for the OMC characterization setup. See the diagram below, but the setup consists of:

  • Faraday isolator/polarization definition
  • HWP+PBS for power splitting into two paths:
    • EOM path
      • Resonant EOM for PDH sideband generation
      • Broadband EOM for frequency scanning
    • AOM path
      • Double-passed ~200-MHz Isomet AOM for subcarrier generation. NOTE: in this case, I have chosen the m = -1 diffraction order due to the space constraints on the table.
  • Recombination of paths on a 50/50 beam splitter---half of the power is lost through the unused port into a black glass dump
  • Coupler for launching dual-field beam into a fiber (to OMC)

OMC_test_setup_IO.jpg 2013-03-27_05.48.20.jpg

Today, we placed some lenses into the setup, in two places:

  1. In the roundabout section of the AOM path that leads to the recombination, to re-match the AOM-path beam to that of the EOM path
  2. After the recombination beam splitter, to match the combined beam mode into the fiber

We (Koji, Lisa, and myself) had significant trouble getting more than ~0.1% coupling through the fiber, and after a while we decided to go to the 40m to get the red-light fiber illuminator to help with the alignment.

Using the illuminator, we realigned the input to the coupler and eventually got much better---but still bad---coupling of ~1.2% (0.12 mW out / 10 mW in). Due to the multi-mode nature of the illuminator beam, the output cannot be used to judge the collimation of the IR beam; it can only be used to verify the alignment of the beam.

With 0.12 mW emerging from the other end of the fiber, we could see the output quite clearly on a card (see photo below). This can tell us about the required input mode. From the looks of it, our beam is actually focused too strongly. We should probably replace the 75mm lens again with a slightly longer one.

Lisa and I concurred that it felt like we had converged to the optimum alignment and polarization, which would mean that the lack of coupling is all from mode mismatch. Since the input mode is well collimated, it seems unlikely that we could be off enough to only get ~1% coupling. One possibility is that the collimator is not well attached to the fiber itself. Since the Rayleigh range within it is very small, any looseness here can be critical.

2013-03-28_02.29.36.jpg

I think there are several people around here who have worked pretty extensively with fibers. So, I propose that we ask them to take a look at what we have done and see if we're doing something totally wrong. There is no reason to reinvent the wheel.

ELOG V3.1.3-